OpenStudy (anonymous):

Write an equation for the line that contains the point (8, -3) and is parallel to the line x + 4y = 9.

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

y=mx+c since the line is parallel to given line m=-1/4

OpenStudy (anonymous):

y=-(1/4)x + c substituting y=-3 and x=8 -3=-2+c c=-1 y=-(1/4)x-1 4y+x+4=0